# Through the River: Understanding the Benefit of Schedule-Free Methods for Language Model Training

### v1

@@ D.1 Proof of Stability Threshold of Schedule-Free Optimizer
Define $(\mathbf{q},a)\coloneqq(\mathbf{q},\lambda_{1}({\mathbf{H}}))$ to be the largest eigenvector/eigenvalue pair of ${\mathbf{H}}$ and $\tilde{x}_{t}=\mathbf{q}^{\top}\mathbf{x}_{t}+\frac{1}{a}\mathbf{q}^{\top}\mathbf{g}$. Then the sequence $\{\mathbf{q}^{\top}{\mathbf{x}}_{t}\}$ diverges if and only if the sequence $\{\tilde{x}_{t}\}$ diverges. By multiplying $\mathbf{q}^{\top}$ on both sides of Equation 1, we get
@@
Since we take $\gamma>0$ and $10\leq\beta<$ 1, the condition $a>\frac{2}{(1-\beta)\gamma}$ implies diverging $\bar{x}_{t}$ as well as $\tilde{x}_{t}$. ∎
D.1.2 Proof of Proposition 4.2
To show Proposition 4.2, we first prove the following reparameterization lemma.
@@
[…]
@@
Let $\{\tilde{{\mathbf{x}}}_{t}\}$ be the iterates of SF-GD on the reparameterized objective $\tilde{g}(\mathbf{w})=g(\mathbf{P}^{-1/2}\mathbf{w})$ with initialization $\tilde{{\mathbf{x}}}_{1}=\mathbf{P}^{1/2}{\mathbf{x}}_{1}$. From Lemma D.1, $\{{\mathbf{x}}_{t}\}$, the iterates of SF-PrecondGDW, satisfy $\tilde{{\mathbf{x}}}_{t}=\mathbf{P}^{1/2}{\mathbf{x}}_{t}$, which implies that if $\tilde{{\mathbf{x}}}_{t}$ diverges then ${\mathbf{x}}_{t}$ also diverges.
From Proposition 4.1, if $\lambda_{1}(\mathbf{P}^{-1}\mathbf{H}+\lambda\mathbf{I})>\frac{2}{(1-\beta)\gamma}$, then $\tilde{{\mathbf{x}}}_{t}$ diverges. This proves the claim. ∎
@@ D.3.2 Deriving the Central Flow of Schedule-Free Scalar Adam
D.4 Omitted Calculations in Section 5
@@
Figure 13 and Figure 14 show that the averaging weights $\{\alpha_{t}\}$ in our refined SF method closely follow the approximation $\alpha_{t}\approx\tfrac{C}{T}(\tfrac{t}{T})^{C-1}$, across different values of $\beta$, $C$, and $T$.

### v4

@@ D.1 Proof of Stability Threshold of Schedule-Free Optimizer
Define $(\mathbf{q},a)\coloneqq(\mathbf{q},\lambda_{1}({\mathbf{H}}))$ to be the largest eigenvector/eigenvalue pair of $\mathbf{H}$} and $\tilde{x}_{t}=\mathbf{q}^{\top}\mathbf{x}_{t}+\frac{1}{a}\mathbf{q}^{\top}\mathbf{g}$. Then the sequence $\{\mathbf{q}^{\top}{\mathbf{x}}_{t}\}$ diverges if and only if the sequence $\{\tilde{x}_{t}\}$ diverges. By multiplying $\mathbf{q}^{\top}$ on both sides of Equation˜1, we get
@@
Since we take $\gamma>0$ and $0\leq\beta<1$, the condition $a>\frac{2}{(1-\beta)\gamma}$ implies diverging $\bar{x}_{t}$ as well as $\tilde{x}_{t}$. ∎
D.1.2 Proof of Proposition˜4.2
To show Proposition˜4.2, we first prove the following reparameterization lemma.
@@
$={\mathbf{P}}^{1/2}{\mathbf{z}}_{t}-\gamma{\mathbf{P}}^{-1/2}\nabla f({\mathbf{y}}_{t})\quad\text{(inductive hypothesis)}$
@@
Let $\{\tilde{{\mathbf{x}}}_{t}\}$ be the iterates of SF-GD on the reparameterized objective $\tilde{g}(\mathbf{w})=g(\mathbf{P}^{-1/2}\mathbf{w})$ with initialization $\tilde{{\mathbf{x}}}_{1}=\mathbf{P}^{1/2}{\mathbf{x}}_{1}$. From Lemma˜D.1, $\{{\mathbf{x}}_{t}\}$, the iterates of SF-PrecondGDW, satisfy $\tilde{{\mathbf{x}}}_{t}=\mathbf{P}^{1/2}{\mathbf{x}}_{t}$, which implies that if $\tilde{{\mathbf{x}}}_{t}$ diverges then ${\mathbf{x}}_{t}$ also diverges.
From Proposition˜4.1, if $\lambda_{1}(\mathbf{P}^{-1}\mathbf{H}+\lambda\mathbf{I})>\frac{2}{(1-\beta)\gamma}$, then $\tilde{{\mathbf{x}}}_{t}$ diverges. This proves the claim. ∎
@@ D.3.2 Deriving the Central Flow of Schedule-Free Scalar Adam
D.4 Omitted Calculations in Section˜5
@@
Figure˜14 and Figure˜15 show that the averaging weights $\{\alpha_{t}\}$ in our refined SF method closely follow the approximation $\alpha_{t}\approx\tfrac{C}{T}(\tfrac{t}{T})^{C-1}$, across different values of $\beta$, $C$, and $T$.